# TrendLSW: Trend and Spectral Estimation of Nonstationary Time Series in R

### v1

@@ 1 Introduction
Two key quantities of interest, the mean function and dependence structure, are amongst the most commonly studied properties in time series analysis. Modelling how these properties evolve is crucial for making informed inference on the data: incorrectly assuming stationary behaviour may lead to drawing misleading conclusions. It is known to be a challenging problem to estimate one of these time-varying properties when the other is stationary, and yet more challenging when both display time-dependent characteristics. Therefore, it is common practice for analysis to be restricted to either the mean or dependence, without a proper consideration of the other.
@@
For modelling nonstationary dependence structures, a number of approaches have been proposed; for an overview, see Dahlhaus, (2012). For example, Priestley, (1965) introduces evolutionary processes, Dahlhaus, (1997) defines the locally stationary Fourier (LSF) processes, whilst Nason et al., (2000) introduce the locally stationary wavelet (LSW) model. Ombao et al., (2002) consider the smoothed localised exponential (SLEX) model, whilst Van Bellegem and Dahlhaus, (2006) discuss nonstationary autoregressive models, where the parameters are allowed to vary with time. Zhou and Wu, (2009) define locally stationary time series as an “input-output” physical system.
Existing software in the area of nonstationary time series predominantly focuses on estimation of either the mean or dependence structure. A comprehensive review of all such packages is outside the scope of this article (in particular any R package implementing a regression approach can be adapted to trend estimation in time series data by using the time index as the covariate). However, to the best of our knowledge, there are no R packages available specifically designed for estimation of trend in the presence of nonstationary dependence. For estimation of nonstationary dependence structure, the R package LSTS (Olea et al.,, 2021) implements the LSF approach. Nason, (2022) implements wavelet-based methods for time series in the wavethresh package, including the LSW model, as well as wavelet-based trend estimation but for a second-order stationary time series. The locits package (Nason,, 2016) allows for time-varying autocovariance estimation under an LSW model, whilst the LSWPlib library (Cardinali and Nason,, 2022) provides locally stationary wavelet packet representations and associated spectral estimation methods. Taylor et al., (2019) extend the LSW approach to multivariate time series in the package mvLSW. Other application-focused packages for nonstationary time series analysis include the RSEIS package (Lees and Harris,, 2023), which has functionality for spectrogram computation and visualisation for seismic data. For trend detection, packages include trend (Pohlert,, 2023) and funtimes (Lyubchich et al.,, 2023). We mention that, for nonparametric regression without a focus on the time series setting, packages include mgcv (Wood,, 2023) and gam (Hastie,, 2023).
@@ 2.1 Wavelets
In the LSW framework (discrete) wavelets act as building blocks in an analogous fashion to Fourier exponentials in the classical Cramér representation for stationary processes. Let $\psi$ be a compactly supported wavelet, for example any within the Daubechies family (Daubechies,, 1992). Denote by $\{h_{k},g_{k}\}$ the low- /high- pass filter pair associated with $\psi$. The pair $\{h_{k},g_{k}\}$ filter a signal into low and high frequency components respectively. Letting $N_{h}$ be the number of non-zero values of $\{h_{k}\}$, the filters are related by the equation $g_{k}=(-1)^{k}h_{N_{h}-k}$ (Nason,, 2008, Equation (2.52)).
@@
For a good introduction of the use of wavelets in statistics, we refer the reader to Antoniadis et al., (2001) and Nason, (2008).
@@ 3.2 First-differenced estimation with modified correction factor
To solve this problem, we can use the commonly used practice of differencing. Before applying the wavelet transform, we first difference the time series to yield $\{\Delta X_{t}=X_{t+1}-X_{t}\}_{t=1}^{n-1}$. Differencing allows us to remove the trend immediately, ensuring that the wavelet transform does not accumulate large bias across increasing scales. Then, the wavelet coefficients $\tilde{d}_{j,k}=\sum_{t}\Delta X_{t}\psi_{j,k}(t)$ are computed on the differenced time series, from which the wavelet periodogram $\tilde{I}_{j,k}=\tilde{d}_{j,k}^{2}$ is constructed.
@@ 4.1 Linear wavelet thresholding estimator
To obtain robust estimation of the trend at the boundary of the series, McGonigle et al., 2022b suggest using a procedure akin to classical boundary handling in wavelet methods, in which a long series of length $4n$ is constructed by reflecting the time series data so that the original series is at the centre of the newly constructed data. Performing trend estimation on this longer series mitigates the fact that, particularly at coarser scales, the wavelet filter can be longer than the available data near the boundary. Note that a nondecimated discrete wavelet transform (NDWT) can be performed instead of the DWT to represent more temporally-localised behaviour in the series, in which case basis averaging is used to obtain the trend estimate after hard thresholding, see e.g., Nason et al., (2000, Chapter 3.12) for more details.
@@
When the NDWT is used, pointwise confidence intervals are computed via the bootstrap, in a similar fashion to the approach of Friedrich et al., (2020). For some chosen total number of bootstrap replications, $B$, bootstrapped replications $\{X_{t}^{(b)}\}_{t=1}^{T}$ for $b=1,\ldots,B$ can be simulated using the spectral estimate and trend estimate, from which bootstrapped trend estimates can be calculated.
@@
$X^{(b)}_{t}=\widehat{T}_{t}+\sum_{j}\sum_{k}\widehat{w}_{j,k}\psi_{j,t-k}\widehat{\xi}_{j,k}$ (12)
are generated, where $\widehat{w}_{j,k}=\widehat{S}_{j}(k/n)^{1/2}$, and $\widehat{\xi}_{j,k}$ are independent, identically distributed standard Normal random variables. For each of the $B$ bootstrapped replications, a bootstrapped trend estimate is obtained using the linear wavelet thresholding procedure as described above. Then, for each time point $t=0,\ldots,n-1$, pointwise $(1-\alpha)$confidence intervals can be calculated using either:

### v2

@@ 1 Introduction
Two key quantities of interest, the mean function and dependence structure, are amongst the most commonly studied properties in time series analysis. Modelling how these properties evolve is crucial for making informed inference on the data: incorrectly assuming stationary behaviour may lead to drawing misleading conclusions. It is known to be a challenging problem to estimate one of these time-varying properties when the other is stationary, and yet more challenging when both display time-dependent characteristics. Therefore, it is common practice for analysis to be restricted to either the mean or dependence.
@@
For modelling nonstationary dependence structures, a number of approaches have been proposed; for an overview, see Dahlhaus, (2012). For example, Priestley, (1965) introduces evolutionary processes, Dahlhaus, (1997) defines the locally stationary Fourier (LSF) processes, whilst Nason et al., (2000) introduce the locally stationary wavelet (LSW) model. Ombao et al., (2002) consider the smoothed localised exponential (SLEX) model, whilst Van Bellegem and Dahlhaus, (2006) discuss nonstationary autoregressive models, where the parameters are allowed to vary with time. Zhou and Wu, (2009) define locally stationary time series as an “input-output” physical system. Other approaches which can capture both time-varying means and time-varying dynamics are time-varying structural autoregression models (e.g. (Kilian and Lütkepohl,, 2017)) and dynamic linear models, see for example, Keele and Kelly, (2006).
Existing software in the area of nonstationary time series predominantly focuses on estimation of either the mean or dependence structure. A comprehensive review of all such packages is outside the scope of this article (in particular any R package implementing a regression approach can be adapted to trend estimation in time series data by using the time index as the covariate). For example, functionality for trend estimation via penalised regression approaches or regression splines (such as in the genlasso (Arnold and Tibshirani,, 2022) and earth (Milborrow,, 2024) packages respectively) could be adapted to serially dependent noise. For estimation of nonstationary dependence structure, the R package LSTS (Olea et al.,, 2021) implements the LSF approach. Nason, (2022) implements wavelet-based methods for time series in the wavethresh package, including the LSW model, as well as wavelet-based trend estimation but for a second-order stationary time series. The locits package (Nason,, 2016) allows for time-varying autocovariance estimation under an LSW model, whilst the LSWPlib library (Cardinali and Nason,, 2022) provides locally stationary wavelet packet representations and associated spectral estimation methods. Taylor et al., (2019) extend the LSW approach to multivariate time series in the package mvLSW. Other application-focused packages for nonstationary time series analysis include the RSEIS package (Lees and Harris,, 2023), which has functionality for spectrogram computation and visualisation for seismic data. For trend detection, packages include trend (Pohlert,, 2023) and funtimes (Lyubchich et al.,, 2023). We mention that, for nonparametric regression without a focus on the time series setting, packages include mgcv (Wood,, 2023) and gam (Hastie,, 2023). However, to the best of our knowledge, there are few R packages available specifically designed for estimation of trend in the presence of nonstationary dependence. Examples include software for dynamic linear models in the dynlm package (Zeileis,, 2019), also available in the nonlinear time series setting in tsDyn (Fabio Di Narzo et al.,, 2009; Stigler,, 2019). However, these are not designed for the locally stationary setting.
@@ 2.1 Wavelets
Broadly speaking, wavelets are localised, oscillatory basis functions that possess several useful properties not typically enjoyed by Fourier trigonometric basis functions. In the LSW framework (discrete) wavelets act as building blocks in an analogous fashion to Fourier exponentials in the classical Cramér representation for stationary processes and spline bases for mean representations. Let $\psi$ be a compactly supported wavelet, for example any within the Daubechies family (Daubechies,, 1992). Denote by $\{h_{k},g_{k}\}$ the low- /high- pass filter pair associated with $\psi$. The pair $\{h_{k},g_{k}\}$ filter a signal into low and high frequency components respectively. Letting $N_{h}$ be the number of non-zero values of $\{h_{k}\}$, the filters are related by the equation $g_{k}=(-1)^{k}h_{N_{h}-k}$ (Nason,, 2008, Equation (2.52)).
@@
For a good introduction of the use of wavelets in statistics, we refer the reader to Antoniadis et al., (2001) and Nason, (2008). Other excellent texts on wavelets include Percival and Walden, (2006) and Vidakovic, (2009).
@@ 3.2 First-differenced estimation with modified correction factor
To solve this problem, we can use the commonly used practice of differencing. Before applying the wavelet transform, we first difference the time series to yield $\{\Delta X_{t}=X_{t+1}-X_{t}\}_{t=1}^{n-1}$. Differencing allows us to remove the trend immediately, ensuring that the wavelet transform does not accumulate large bias across increasing scales. Then, the wavelet coefficients $\tilde{d}_{j,k}=\sum_{t}\Delta X_{t}\psi_{j,k}(t)$ are computed on the differenced time series, from which the wavelet periodogram $\tilde{I}_{j,k}=\tilde{d}_{j,k}^{2}$ is constructed. It is possible to use higher order differencing, e.g. second differencing, however we advise against this due to the inherent loss of information and poorer empirical performance as demonstrated in McGonigle et al., 2022a .
@@ 4.1 Linear wavelet thresholding estimator
To obtain robust estimation of the trend at the boundary of the series, McGonigle et al., 2022b suggest using a procedure akin to classical boundary handling in wavelet methods, in which a long series of length $3n$ is constructed by reflecting the time series data so that the original series is at the centre of the newly constructed data. Performing trend estimation on this longer series mitigates the fact that, particularly at coarser scales, the wavelet filter can be longer than the available data near the boundary. Note that a nondecimated discrete wavelet transform (NDWT) can be performed instead of the DWT to represent more temporally-localised behaviour in the series, in which case basis averaging is used to obtain the trend estimate after hard thresholding, see e.g., Nason et al., (2000, Chapter 3.12) for more details. The key difference between the DWT and NDWT is that DWT forms an orthogonal basis whereas the NDWT is an overcomplete basis representation. The DWT downsamples the data at each step of the algorithm to obtain a sparse (efficient) representation of the underlying signal. The disadvantage of this is that if the data is shifted by a single place to the left or right, the DWT representation can change dramatically (not just shifted one to the left/right). In contrast, the NWDT is translation-invariant as it considers all potential paths of shifting and downsampling the data at each step of the algorithm and does not suffer from this drawback.
@@
When the NDWT is used, pointwise confidence intervals are computed via the bootstrap, in a similar fashion to the approach of Friedrich et al., (2020). For some chosen total number of bootstrap replications, $B$, bootstrapped replications $\{X_{t}^{(b)}\}_{t=0}^{n-1}$ for $b=1,\ldots,B$ can be simulated using the spectral estimate and trend estimate, from which bootstrapped trend estimates can be calculated.
@@
$X^{(b)}_{t}=\widehat{T}_{t}+\sum_{j}\sum_{k}\widehat{w}_{j,k}\psi_{j,t-k}\xi^{{\color[rgb]{0,0,0}(b)}}_{j,k}$ (12)
are generated, where $\widehat{w}_{j,k}=\widehat{S}_{j}(k/n)^{1/2}$, and $\xi^{{\color[rgb]{0,0,0}(b)}}_{j,k}$ are independent, identically distributed standard Normal random variables. For each of the $B$ bootstrapped replications, a bootstrapped trend estimate $\{\widehat{T}_{t}^{(b)}\}_{t=0}^{n-1}$ is obtained using the linear wavelet thresholding procedure as described above. Then, for each time point $t=0,\ldots,n-1$, pointwise $(1-\alpha)$confidence intervals can be calculated using either:
@@ 6.1 Bioluminescence of Caenorhabditis elegans
It can be beneficial to fit a range of wavelets and assess the visual fit to the data in order to choose the most appropriate wavelet for trend estimation. In the code below, we compute the trend estimate for a range of wavelets.
R> out.EP1 <- TLSW(celegensbio, T.filter.number = 1,                    T.family = "DaubExPhase", T.transform = "nondec")R> out.EP4 <- TLSW(celegensbio, T.filter.number = 4,                   T.family = "DaubExPhase", T.transform = "nondec")R> out.LA10 <- TLSW(celegensbio, T.filter.number = 10,                    T.family = "DaubLeAsymm", T.transform = "nondec")
We can then visually compare the fitted trends, as shown in Figure 11.
R> plot(out, plot.type = "trend", trend.plot.args = list(        ylab = "Luminescence (AU)",        T.col = "blue", T.lwd = 2, main = ""),        cex.lab = 1.3, col = "grey60")R> lines(out.EP1\$trend.est\$T, col = "black", lwd = 2, lty = 2)R> lines(out.EP4\$trend.est\$T, col = "red", lty = 3, lwd = 2)R> lines(out.LA10\$trend.est\$T, col = "green", lty = 4, lwd = 2)R> legend(10, 85000, c("EP10", "EP1", "EP4", "LA10"),       lty = c(1, 2, 3, 4),lwd = c(2, 2, 2, 2),       col = c("blue", "black", "red", "green"), bty = "n", cex = 1.2)
We compare the fits produced by the Daubechies Extremal Phase (EP) wavelets with 1, 4, and 10 vanishing moments, as well as the Daubechies Least Asymmetric (LA) Wavelet with 10 vanishing moments. The fits for the EP1 and EP4 wavelets appear to be noticeably worse than those for the EP10 and LA10, which show better adaptivity to the data and are almost identical. Heuristically, we can also compute the root mean squared error (RMSE) of the trend estimates, which identify the EP10 and LA10 wavelets as being indistinguishable (RMSEs of $3682.85$ and $3682.42$ respectively), and superior to the EP1 and EP4 wavelets (RMSEs of $4062.61$ and $3990.25$ respectively). The code for this is given below.
R> n <- length(celegensbio)R> trends <- list(EP1 = out.EP1\$trend.est\$T,EP4 = out.EP4\$trend.est\$T,               EP10 = out\$trend.est\$T, LA10 = out.LA10\$trend.est\$T)R> trends.rmse <- lapply(trends,                         function(y){sqrt(sum((y - celegensbio) ^ 2) / n)})
Figure 11: Comparison of TLSW trend fits obtained using: the EP10 wavelet (blue solid line), EP1 wavelet (black dashed line), EP4 wavelet (red dotted line), LA10 wavelet (green dot-dashed line).
@@ 7 Concluding remarks
Future development of TrendLSW will include adding further functionality including a predict method for TLSW objects.